OpenStudy (anonymous):

on a survey of 550 people, 62 percent like tasty pizza. estimate how many people like tasty pizza.

OpenStudy (anonymous):

\[\frac{ 62 }{ 100 }\times550\]

OpenStudy (dan815):

some number of people like pizza out of 550 which is 62 of them so x/550=62%

OpenStudy (dan815):

x/550=62/100 solve for x again

OpenStudy (anonymous):

\[=341\] Therefore 342 people like tasty cheese

OpenStudy (dan815):

which is 62%* of them

OpenStudy (anonymous):

*431 people like tasty cheese

OpenStudy (dan815):

no dian u had it right the first time i think

OpenStudy (anonymous):

Ya

OpenStudy (dan815):

when ever ur doing percentage questions always think fractions, theres gonna be fractions popping up all over

OpenStudy (anonymous):

Wait no because my answers are: A:330 B:25 C:250 D:35

OpenStudy (dan815):

well if u understand the logic of how you get the answer, there shud be no doubt in your mind if its right or wrong! so u tell me

OpenStudy (anonymous):

I don't know!!! Remember..I suck at math!?!?!? Haha

OpenStudy (dan815):

no u dont xD math is just a language u just gotta learn it, its like learning english

OpenStudy (anonymous):

just relook at your question and follow the same approach, you'll get the answer!

OpenStudy (anonymous):

I'm trying! :(

OpenStudy (dan815):

maybe u meant 60%?

OpenStudy (anonymous):

^ that would give you a correct number

OpenStudy (anonymous):

Ya but it says on my paper 62

OpenStudy (anonymous):

330?

OpenStudy (anonymous):

yes

OpenStudy (dan815):

the answer is none of the above